[Brahms-l] Shift report 20050521 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sat May 21 2005 - 23:55:51 EDT
Shifter leader was kh
Shifters were 


Had beam during a good amount of time (6 hours), but most of the data taken during this shift was with D1 off because D1 tripped and aborted the beam.  People did not like that and eventually they asked us (right before this shift started) to run without D1 until the problem was fixed.  The problem was diagnosed to be that the temperature in the PS room was too high.  So support came and put freon into the AC and the room cooled.  At 23:00, we were given the go-ahead to turn D1 back on sice the room was cool enough.

The runs starting from 14823 are the only ones in this shift that have D1 on.

No detector problems were obvious.  In particular, CC counters continued to work.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat May 21 23:56:05 2005

This archive was generated by hypermail 2.1.8 : Sat May 21 2005 - 23:56:18 EDT